IMACS progress report no. 48
B. Bigelow 5 April 2002
I) Project Status
Currently scheduled tasks: Status:
a) DSS redesign (Hare) IP
b) DSS drafting (Cleary) IP
c) DSS fabrication (Kowal, Castillo) IP
d) Calibration system mech. des. (Hare) IP
e) Long camera opto-mech assembly (Bigelow) IP
f) Field lens integration & assy (Hare) on sched.
g) Short camera mechanics fab (M&T) done
h) Principal Guider Camera testing IP
i) Shack-Hartmann Guider Camera testing IP
j) Science Array controller fab (Burley) 95% compl.
k) CCD array assembly (Thompson/Burley) on sched.
l) Filter procurement (Sutin) IP
m) Grism procurement (Sutin) IP
n) Motion control electronics design (Bagish) done
o) Instrument electronics assembly (Bagish) on sched.
p) MC fab. (Asa) 70% compl.
q) Control software design (Birk) on sched.
r) Instrument structure fabrication (M&T) done
s) Instrument enclosures design (Gunnels) on sched.
t) Mask cutter fabrication (Convergent) done
Project events in March:
- The instrument structure was completed, delivered and assembled.
- The short camera cells, barrels, and spacers were delivered by M&T.
- The slit mask cutting laser was completed, tested, and shipped to LCO.
The laser can cut features down to 50 microns width, or 0.15 arcsec
on the slit masks. Our internal slit mask cutting software was
successfully tested as well.
- Collimator assembly was completed, and the collimator is now mounted
on the structure. Element spacings were correct at assembly, and no
changes to spacers were required.
- CCD array assembly is progressing, and flatness of the array will be
limited by the individual chip flatness, to about 15 microns. Pixel
alignment of the chips (end-to-end, side-to-side) is to better than
1 pixel over 8192 pixels.
Latest project images:
Final Assy: http://www.ociw.edu/instrumentation/imacs/devel/final/index.html
Collim. assy: http://www.ociw.edu/instrumentation/imacs/devel/coll/index.html
II) Optics (Sutin)
1) Short camera spherical lenses: Everything finished except S07.
S07 is not CaF2, so it should not run over too long. I don't
yet have an expected time of completion.
2) Short camera aspheres: Tinsley is about a month late. They
finished grinding S08 and will get a 20% progress payment.
3) STF gave us a coating bid for the Short camera and grisms which
is just under our budget. We are taking it, and will start
4) We have material for the 600 l/mm grism, but no manufacturer.
Zygo no-bid for doing a contact-bond, but I may convince them
to do a glue-bond.
5) Barr promises to send the R filters by the end of April.
6) I asked Custom Scientific to bid on the Sloan filters. They
will be cheaper, but cannot test homogeneity.
7) Schott has acknowledged that I wrote them a letter about CaF2,
but that is it.
8) The DI mirror was send to TORC and re-polished. It is now being
send to Denton for enhanced Aluminum. Silver usually lasts about
ten years, and that is too short for a potted mirror. Since we
need one now for testing, I rented a replacement from Davidson
Optronics for a month.
9) If desired, we can purchase a second DI mirror for about $1400
for the substrate plus $550 for either enhanced Aluminum or Silver
10) Max Levy came in at about $2800 to make the CF guider mask, but
it now turns out that our laser should be able to do the job
in-house. That is the current plan.
III) Electronics (Bagish)
- Motion Control Electronics Chassis #3 99% complete. Waiting on one jack to
come in to finish.
- Motion Control Electronics Chassis #4 is in process of being wired up.
Expect completion by end of month. Short camera detector focus driver was
moved from Chassis 2 to join Long camera detector focus in Chassis #4 so
that both are in the same chassis. Chassis 4 also controls IMACS structure
- Front panels have been machined. Front panels need to be grained, anodized
- Rack side panel needs openings to be cut for mounting Rittal heat-exchanger.
- Ball-bearing slides have been modified to fit racks. Four sets have been
installed in Rack #1 (Carnegie-built electronics).
- Rack # 2 came in. Off-the-Shelf electronics have been installed for fit in
- Cables assembly continues.
- Schematics continue to be updated with corrections and fixes.
IV) Software (Birk/Clardy)
We received the 15:1 gears and installed new drive-chains. The motor
can now easily move the guider against gravity.
Next: check for any granularity in the theta-axis.
Question: Do we need a closed-loop control of the theta-axis when
IMACS rotates? The answer will have to wait until the guiders are
mounted on the instrument structure.
- Science array software:
The sub-raster readout works in simulation mode. The user can define
sub-raster geometry by a) using a dialog-box, or b) point+click on
a previously taken image on the quick-look tool.
Next: I hope to test array readout later in April.
Clardy: (Bigelow for Clardy)
- Mask cutting software was successfully tested in March at Convergent.
- Additional software for cutting LDSS2 masks with the IMACS laser will
V) Detector Systems (Thomson/Burley)
- Finished final profiling of all 11 CCD's, peak to valley range for
all 9 chips (three points per chip) is 7.5 microns.
- Aligned 6 of 11 CCDs. These are typically parallel to plus/minus 0.5
pixels, with end to end separations constant to 1 pixel.
- Assembled temperature sensing/heating fixtures for the 8 CCD's in
- Final assembly of MIKE dewar hardware. The red camera will be assembled
and tested within a week.
- Six more guider cameras have been built and need testing. A seventh is
waiting on a thermoelectric cooler. Three more CCD's have been
ordered, this will bring us to a total of 23 cameras.
- Tested two CCD's in a dewar, configured as a 4K x 4K array. No cross
talk was evident in flat field tests. In early April we will test
measure the cross talk on a preamp card using the spectrum analyzer.
- Tested new DSP board -- one remaining problem with optical module
footprint (barnacled for now).
- Tested optical link -- partially successful, some EPLD reprogramming
- Assemble IMACS dewar with platten and boats and temperature probes. Test
cold, check out thermal control.
- Balance dewar translation stage.
- Assemble CCD's onto platten and profile warm.
- Cold test of array likely to occur in early May.
VI) Mechanics (Bigelow/Hare)
1) Accepted and installed the instrument structure at SBS
2) Accepted, tested the slit mask cutting laser at Convergent, Sturbridge, MA
3) Checked lens spacings in collimator - OK - installed in structure
4) Check and adjust lens spacings in long camera - started with L01
5) Mask cutting tooling designed, built, and tested with laser
6) Short camera mechanics completed and delivered
1) Final assembly and spacing measurements for the long camera
2) Attempt main optics assembly into structure (field lens, collimator, lcam)
3) Final assembly and alignment of SMS on structure
4) Design of cable wrap and instrument enclosure details (instrument dark slide)
5) Design/fab dewar dummy mass for flexure testing cameras
6) Design/fab CCD array dummy mass for flexure testing dewar
* Revising drive mechanism on Principle and Shack-Hartmann guide cameras to
resolve flexure issues.
* Continued opto-mechanical and electro-mechanical assembly and testing of the
Principle and Shack-Hartmann Guide Cameras.
* Fabrication of Field Lens Tooling Assembly complete.
* Fabrication of the Disperser Server Wheel Assembly and Main DSS Support
* Design of limit switch hardware for the Grating Tilt Mechanism is complete.
Part drafting underway.
* Continue assembly, testing, and rework of all guide cameras as necessary.
* Mount the Field Lens in its cell.
* Test the Clamping Assembly.
* Assemble and test one Grating Tilt Mechanism.
* Revise and fabricate DSS Drive Assembly
* Assembly and testing of the Disperser Server System.
* Fabrication of the second shutter.
* Resume design work of the Calibration Unit.
VII) Fabrication (Kowal)
1) Completed Laser Tooling.
2) Completed Field Lens Tooling.
3) Mask Shuttle Support was finished.
4) Electronic Panels & chassis are done.
5) Convergent Prima laser training.
Expectations for April:
1) Dewar Front Plate
2) Tapered Guide Camera Bases
3) Covers for Collimator and Field Lens
4) Modification of various parts as needed
VIII) Other Business
Next meeting: 10:15, 5/8/02